# Building on Efficient Foundations: Effectively Training LLMs with Structured Feedforward Layers

### v1

@@ Abstract
State-of-the-art results in large language models (LLMs) often rely on scale, which becomes computationally expensive. This has sparked a research agenda to reduce these models’ parameter count and computational costs without significantly impacting their performance. Our study focuses on transformer-based LLMs, specifically targeting the computationally intensive feedforward networks (FFN), which are less studied than attention blocks. We consider three candidate linear layer approximations in the FFN by combining efficient low-rank and block-diagonal matrices. In contrast to many previous works that examined these approximations, our study i) explores these structures from the training-from-scratch perspective, ii) scales up to 1.3B parameters, and iii) is conducted within recent Transformer-based LLMs rather than convolutional architectures. We first demonstrate they can lead to actual computational gains in various scenarios, including online decoding when using a pre-merge technique. Additionally, we propose a novel training regime, called self-guided training, aimed at improving the poor training dynamics that these approximations exhibit when used from initialization. Experiments on the large RefinedWeb dataset show that our methods are both efficient and effective for training and inference. Interestingly, these structured FFNs exhibit steeper scaling curves than the original models. Further applying self-guided training to the structured matrices with 32% FFN parameters and 2.5$\times$ speed-up enables only a 0.4 perplexity increase under the same training FLOPs. Finally, we develop the wide and structured networks surpassing the current medium-sized and large-sized Transformer in perplexity and throughput performance. Our code is available at https://github.com/CLAIRE-Labo/StructuredFFN/tree/main.
(a) Scaling curves
(b) Performance comparisons
Figure 1: (a) The scaling curves during training between the standard Transformer and the modified version with a Structured FFN, which retain 63% and 32% of the original parameters, respectively. The loss curve of the structured FFN is steeper, with a larger gap compared to the dense one at the beginning and a much smaller gap at larger sizes. (b) Performance between dense Transformer, Structured FFN with 32% parameters either trained with self-guided training or more tokens across four sizes. Circle size indicates model FLOPs. To enable straightforward comparison, we controlled their training FLOPs to be the same.
@@ 1 Introduction
Structured linear transformations, such as low-rank or block-diagonal matrices, are important paradigms for reducing the computational cost of feedforward layers. However, they have not yet been thoroughly explored at a sufficient scale to reduce pre-training costs and latency incurred during the inference phase in modern LLM architectures, where the main focus so far has been on improving the efficiency of the self-attention mechanism.
@@
In this work, we investigate structured matrices for FFN blocks from the train-from-scratch aspect, first identifying their efficiency and optimization challenges and then presenting experimental results, analyzing and characterizing the behavior of models trained with structured matrices, and comparing their results. We consider three efficient linear parametrizations: LowRank, BlockShuffle (comprising two block-diagonal matrices), and BlockDense (a combination of dense and block-diagonal matrices). First, while they have demonstrated materialized computational gains, they face challenges in the practical online decoding scenario of LLM, which may process only limited input tokens at one time, leading to under-utilization of computing resources and decreased efficiency due to the additional linear projection. We address this with a pre-merge technique that restores efficiency to the original dense parametrization. Second, regarding training, we noticed that these parametrizations of the FFN blocks would suffer from either slower convergence or less stable training dynamics. To counter this, we propose a flexible and fast method we refer to as self-guided training. It employs a dense matrix as a residual component during the initial training phase, steering the training process away from suboptimal starting points gradually.
We conduct our experiments at scale on Transformers ranging from 110M to 1.3B parameters. By replacing the original FFN with efficient alternatives that use only 32% of the FFN parameters, we boost the training speed of the 1.3B model by 1.35$\times$ while only increasing the perplexity by 1 point. Interestingly, scaling curves in 1(a) indicate that structured matrices yield steeper loss scaling curves than the traditional Transformer at its optimal trade-off, suggesting a high potential for even better performance at larger scales. Furthermore, our results in 1(b) demonstrate that self-guided training enhances learning dynamics and thus the final performance, such as the 0.4 perplexity loss on the 1.3B model with the same training FLOPs and 2.5$\times$ faster FFN with 32% original parameters at inference.
Based on these findings, we design wide and structured networks at medium and large sizes using the efficient GQA [8] technique for attention and structured linear parameterizations for FFN blocks. With the same training FLOPs, our model outperforms the existing Transformer with GQA, boosting maximum throughput by 8% and 17%, respectively, while achieving slightly lower perplexity. As the first work to explore structured matrices at the scale of recent LLMs, we hope our findings and results will shed new light on the study of efficient NLP architectures.
@@ 2 Method
Multiple techniques have been proposed to approximate linear projections from sparsity to low-rank approximations. We provide an in-depth discussion of existing literature on approximating linear layers in section 3, that better contextualizes our work. We focus on structured approximations of linear projections $g({\bm{x}})={\bm{W}}{\bm{x}}$ that have the form $g({\bm{x}})={\bm{U}}({\bm{V}}{\bm{x}})$, where ${\bm{U}}$ and ${\bm{V}}$ are structured matrices, e.g. low-rank or block-diagonal. We opt for this particular form because it allows us to readily exploit the computational gains on existing hardware using existing libraries with minimal alteration. Such approximations have been previously studied in different contexts. Our contributions are exploring them (i) to approximate FFN layers of transformers, (ii) when applied from initialization, (iii) testing them at up to 1.3B scale, and providing scaling curve studies.
@@ 2.1 Structured linear parametrization
We explore three structures to approximate the standard linear layer ${\bm{W}}{\bm{x}}$, maintaining its input dimension $N$ and output dimension $M$ of weight ${\bm{W}}$.
@@
The previous parametrization incorporates two additional shuffle operations, which can be slow on the device. We propose an alternative efficient parametrization that retains the first block-diagonal matrix ${\bm{V}}^{b}$ but substitutes the second with a dense matrix ${\bm{U}}^{r}$. Thus, we can mix the features between blocks without permuting the inner features.
The formula is defined as ${\bm{W}}{\bm{x}}\approx{\bm{U}}^{r}({\bm{V}}^{b}{\bm{x}})$, where ${\bm{V}}^{b}$ is the block-diagonal matrix with $b$ blocks in shape $\frac{R}{B}\times\frac{N}{B}$, and ${\bm{U}}^{r}\in\mathbb{R}^{M\times R}$. Technically, the second projection does not need to be a low-rank approximation because $R$ can be larger than $M$. Nevertheless, in practice, we chose $R<M$ to limit the search space of this work, and thus use the superscript $r$ for the second matrix. The parameters of this method are determined by two variables $B$ and $R$, cutting the original burden from $M\cdot N$ to $R\cdot(M+\frac{N}{B})$. Note that BlockDense can recover the LowRank approximation if we set $B=1$ and $R<\min(M,N)$.
Implementation. We take Dao et al. [12]’s implementation for the BlockShuffle and carefully manage memory copies for BlockDense. In our experiments, we chose $B$ as a common divisor of $M$ and $N$ or $R$. Proper initialization is also investigated in Sec. B.1.
We limit our exploration of efficient linear layers within the FFN blocks.These typically have $8\cdot H^{2}$ parameters and MAC, where $H$ standards for hidden state dimension and $4\cdot H$ as the intermediate hidden size of FFN. In contrast, the proposed parametrizations have:
@@ 2.3 Addressing the optimization challenge
Using the efficient parametrization from initialization can suffer from optimization difficulty because the deep linear parametrization introduces additional symmetries [^2] Symmetries arise because such a factorization ${\bm{U}}{\bm{V}}$ is not unique; for any invertible matrix ${\bm{C}}$ of size $R\times R$ we have ${\bm{U}}{\bm{V}}={\bm{U}}{\bm{C}}{\bm{C}}^{-1}{\bm{V}}=({\bm{U}}{\bm{C}})({\bm{C}}^{-1}{\bm{V}})=\tilde{{\bm{U}}}\tilde{{\bm{V}}}$. BlockShuffle can also be included by fusing the shuffle operation into block-diagonal matrices. , which is a source of proliferation of saddle points and generally less smooth loss function as pointed out in [17, 18]. We hypothesize that this makes poorer learning dynamics of the structured parametrization. Empirically, we show that the deep linear form of ${\bm{U}}({\bm{V}}{\bm{x}})$ in Fig. 3 leads to instability and loss spike or to slow convergence compared to the dense linear projection. We further elaborate on this, highlighting how the inconsistency of gradient updates between the structured parametrization and original linear projection affect learning dynamics in Appendix A.2.
@@
(b) Slower convergence
Figure 3: Poor training dynamics: Training dynamics of LowRank with rank of 128 under different training configurations. Curves correspond to a 4-layer Transformer with a model width of 768 on WikiText-103. We apply self-guided training in the first half of training. Refer to Sec. A.1 for additional details.
@@
To motivate our proposal, we start by finding that the updates on ${\bm{U}}{\bm{V}}$ scales the function of the backpropagated gradients ${\bm{g}}$ (see App. A.2), then turn into the typical training dynamics with gradients. An issue that needs to be addressed in the early stage of training is feature specialization when the learning process assigns semantics to the different hidden units of the model, sometimes also phrased as identifying the winning ticket [19]. In this process, certain weights will need to be suppressed, and symmetries in the parametrization must be resolved.
@@ 3 Related work
Efficient techniques for training LLMs Recent advancements in attention mechanisms have significant improvements in the efficiency of attention [20, 21, 22, 8, 23, 24, 25], and the focus has shifted towards improving FFNs in LLMs, which contributes to at least half of the training time. Dynamic architectures such as mixtures of experts [26, 27, 28], or optimizers with faster convergence [29, 30] have been popular in improving training efficiency. Moreover, Xia et al. [31] employs structured pruning with learned sparsity masks and a dedicated data-loading policy to reduce the training budget.
@@ 4 Experiments
In our experiments, we analyzed the scaling curves, efficiency, and training challenges of structured parametrization of FFN module in LLMs.
Model We use a basic Transformer architecture [41, 2] with Rotary Embedding [42] and SentencePiece Tokenizer [43]. The original FFN module includes two linear layers and a GeLU activation. Our model sizes include 110M (Transformer-s), 335M (Transformer-m), 729M (Transformer-l) and 1.3B (Transformer-xl) parameters, with sequence lengths of 1024 and model widths from 768 to 2048. We only replace FFN modules with efficient linear layers, as the attention module is well-studied [8, 22]. We explore two alternatives to dense FFNs, reducing parameters to 63% or 32% of the original size by adjusting rank and block number. The first FFN module is not replaced to avoid non-negligible performance degradation.
@@
Training We use the RefinedWeb dataset [44] and randomly select 0.5B tokens for validation, reserving the rest for training. Following scaling laws [45], we allocate tokens at 20 times the number of each baseline model’s parameters. We train using A100 80G GPUs with mixed precision (bfloat16 and float32) and adopt fast CUDA kernels like Flash Attention [24] in all experiments for fair comparisons. Hyperparameters, including learning rates and the global number of tokens per batch, are set according to recent studies [46, 47]. We measure training FLOPs as in Megatron [48], including all matrix multiplications. For more details on our experimental settings, see Appendix C.
Figure 4: Latency and FFN sizes: Latency over 30K tokens of different FFN sizes. The intermediate size of the FFN is set to be 4 times the FFN width. In (a), we show latency results of 63% and 32% FFN parameters across different FFN widths. (b) shows the latency of LowRank, BlockShuffle, and BlockDense with different numbers of parameters by varying the rank and number of blocks from the dense FFN with a width of 5120.
Table 1: Performance comparisons of different FFN parameterizations: Performance of two sizes of different structured matrices with 63% and 32% of the original FFN module’s parameters. We report model size, total FFN size, training tokens, training FLOPs, and training time. Note that the total structured FFN is not exactly 63% of the original because we don’t replace the first FFN module. Also, the BlockDense is slightly smaller for -m and -xl models to ensure the rank is a multiple of 256 when matching parameters. Loss and perplexity (PPL) are evaluated on a 0.5B token validation set. Please refer to Table 9 for the whole table.
Architecture Model FFN Training PPL Size (M) Size (M) Tokens (B) FLOPs Time (h) Transformer-xl 1274 805 25.5 2.10e+20 352.2 12.46 63% LowRank (R=1024) 985 516 25.5 1.66e+20 302.2 12.86 BlockDense (B=4, R=1536) 955 486 25.5 1.61e+20 298.7 12.97 BlockShuffle (B=2) 985 516 25.5 1.66e+20 330.6 12.98 32% LowRank (R=512) 744 275 25.5 1.29e+20 260.2 13.55 BlockDense (B=4, R=768) 728 260 25.5 1.27e+20 261.2 13.74 BlockShuffle (B=4) 744 275 25.5 1.29e+20 284.9 13.81
4.1 Evaluating latency results
Large number of tokens Using large $T$, the standard linear layers and our efficient structured parametrizations become computation-bound where FLOPs become a latency bottleneck [15]. This setting mainly concerns training, the prefill phase of inference, and extensive offline decoding. By reducing FLOPs, it can be seen that in 4(a) and 4(b), our designs can improve the latency consistently. In 4(b), we showed the acceleration of structured parametrization for an FFN module across different ranks or blocks. For BlockDense, the block number does not affect latency performance, all achieving a 1.72x speed-up with 50% fewer parameters. 4(a) compares FFN and different structured FFN results with 63% and 32% parameters across different widths. GPU utilization improves with increasing width. From width 1536, LowRank and BlockDense with 63% FFN parameters achieve a 1.4$\times$ speed-up and a 2.5$\times$ speed-up. BlockShuffle offers modest improvements, with 1.1$\times$ and 2.0$\times$ speed-ups for the two cases, respectively.
Small number of tokens FFN can be parallelism-bound with small $T$ on the A100 GPUs, for example, when $T=16$. With increasing $T$, FFN becomes memory-bound and will eventually be computation-bound. Small $T$ can occur during online and offline decoding stages when unrolling the model step by step. As discussed earlier, our pre-merge method can alleviate the parallelism-bound issue and maintain the same latency with dense matrices. Fig. 5 shows the latency results for three different widths, varying the batch of tokens to determine when to use efficient alternatives or choose pre-merged dense matrices. For example, with a 2048-width FFN, it is difficult to fully utilize resources on GPU with limited tokens. The performance improves significantly when using width 5120 and 6144, such as speed improvements of 2.63$\times$ speed-up of LowRank with 32% FFN parameters on $T=2048$ and 2.81$\times$ acceleration of BlockDense with 32% parameters on $T=1536$.
4.2 Findings on efficient training

### v2

@@ Abstract
State-of-the-art results in large language models (LLMs) often rely on scale, which becomes computationally expensive. This has sparked a research agenda to reduce these models’ parameter counts and computational costs without significantly impacting their performance. Our study focuses on transformer-based LLMs, specifically targeting the computationally intensive feedforward networks (FFNs), which are less studied than attention blocks. We consider three structured linear parameterizations of the FFN using efficient low-rank and block-diagonal matrices. In contrast to many previous works that examined these approximations, our study i) explores these structures from a training-from-scratch perspective, ii) scales up to 1.3B parameters, and iii) is conducted within recent Transformer-based LLMs rather than convolutional architectures. We demonstrate that these structures can lead to actual computational gains in various scenarios, including online decoding when using a pre-merge technique. Additionally, we propose a novel training regime, called self-guided training, aimed at improving the poor training dynamics that these approximations exhibit when used from initialization. Interestingly, the scaling performance of structured matrices is explored, revealing steeper curves in scaling training FLOPs, along with a favorable scaling trend in the overtraining regime. Specifically, we show that wide and structured networks can utilize training FLOPs more efficiently, with fewer parameters and lower loss than dense models at their optimal trade-off. Our code is available at https://github.com/CLAIRE-Labo/StructuredFFN/tree/main.
Table 1: Better training FLOPs utilization of the wide and structured Networks: we compare dense Transformers trained according to their optimal scaling law [1], efficient Transformers (GQA) [2] with high throughput, and our wide and structured networks using LowRank parameterization in the FFN module and reduced attention heads, under the same training FLOPs. TP (throughput) refers to the maximum throughput measured over a generation length of 256.
[…]
Figure 1: Steeper scaling curves of LowRank with 63% or 32% FFN parameters. For more results, see Sec. 4.2.
@@ 1 Introduction
Structured linear transformations, such as low-rank or block-diagonal matrices, are important paradigms for reducing the computational cost of feedforward layers. However, they have not yet been thoroughly explored at a sufficient scale to reduce pre-training costs and latency of the inference phase in modern LLM architectures, where the main focus so far has been on improving the efficiency of the self-attention mechanism.
@@
In this work, we investigate structured matrices for FFN blocks from the train-from-scratch aspect, first identifying their efficiency and optimization challenges and then presenting experimental results, analyzing and characterizing the behavior of models trained with structured matrices, and comparing their results. We consider three efficient linear parametrizations: LowRank, BlockShuffle (comprising two block-diagonal matrices), and BlockDense (a combination of dense and block-diagonal matrices). First, while they have demonstrated materialized computational gains, they face challenges in the practical online decoding scenario of LLM, which may process only limited input tokens at one time, leading to under-utilization of computing resources and decreased efficiency due to the additional linear projection. We address this with a pre-merge technique that restores efficiency to the original dense parametrization. Second, we observe that these parameterizations of the FFN blocks are harder to train than standard linear layers, often exhibiting poorer training dynamics like loss spikes. To counter this, we propose a flexible and fast method we refer to as self-guided training. It employs a dense matrix as a residual component during the initial training phase, steering the training process away from suboptimal starting points gradually.
We conduct our experiments at scale on Transformers ranging from 110M to 1.3B parameters by replacing the traditional heavy FFN with structured matrices. Our experiments first show the scaling behavior of these structured linear parameterizations and then illustrate how our proposed methods address their general efficiency and optimization challenges. First, we examine scaling performance from the perspectives of training compute and model size, highlighting the potential of structured matrices. By controlling for the same training FLOPs, we find that structured FFNs show steeper loss scaling curves than traditional Transformers at optimal trade-offs (See Fig. 4 and Fig. 1). Specifically, as seen in Table 1, our wide and structured networks use training FLOPs more efficiently, needing fewer parameters (464M vs. 729M) and achieving a 17% throughput boost on the -l scale, while still maintaining slightly better perplexity compared to the efficient Transformer [2]. Beyond training compute scaling, we also scale model size in the overtraining regime, with Fig. 5 showing favorable scaling trends for our wide and structured models. Second, our results on efficiency show that structured FFNs, with only 32% of the FFN parameters, can boost the training speed of the 1.3B model by 1.35$\times$. Furthermore, self-guided training enhances the performance of all three structured matrices (e.g., reducing the perplexity gap of LowRank to about 0.4) without affecting the inference time speed-up.
As the first work to explore structured matrices at the scale of recent LLMs, we hope our findings and results will shed new light on the study of efficient NLP architectures. Our contributions can be categorized into the following three aspects:
  1. 1.
    We investigate three types of structured matrices in Transformer pretraining and demonstrate their favorable scaling behavior compared to dense models. This is revealed through the study of scaling laws for training FLOPs, as well as model size scaling in the overtraining regime, showing that wide and structured networks can be strong candidates for architecture design.
  2. 2.
    We conduct an efficiency study of these structured matrices across various scenarios. We propose a pre-merge technique to maintain speed in a specific case and show the effective speed-up performance of structured matrices in other scenarios.
  3. 3.
    We identify optimization challenges in structured matrices and introduce a method called self-guided training, which efficiently improves training dynamics and boosts the final performance for all three types of structured matrices.
@@ 2 Method
Multiple techniques have been proposed to approximate linear projections from sparsity to low-rank approximations. We provide an in-depth discussion of existing literature on approximating linear layers in section 3, that better contextualizes our work. We focus on structured approximations of linear projections $g({\bm{x}})={\bm{W}}{\bm{x}}$ that have the form $g({\bm{x}})={\bm{U}}({\bm{V}}{\bm{x}})$, where ${\bm{U}}$ and ${\bm{V}}$ are structured matrices, e.g. low-rank or block-diagonal. We opt for this particular form because it allows us to readily exploit the computational gains on existing hardware using existing libraries with minimal alteration. Such approximations have been previously studied in different contexts. Our contributions are exploring them (i) to approximate FFN layers of transformers, (ii) when applied from initialization, (iii) testing them at up to 1.3B scale, investigating their general bottlenecks and providing scaling analyses.
@@ 2.1 Structured linear parametrization
We explore three structured matrices to approximate the standard linear layer ${\bm{W}}{\bm{x}}$, maintaining its input dimension $N$ and output dimension $M$ of weight ${\bm{W}}$.
@@
The previous parametrization incorporates additional shuffle operations, which can be slow on the device. We propose a natural intermediate candidate between LowRank and BlockShuffle, combining the block-diagonal projection ${\bm{V}}^{b}$ with a dense or low-rank projection ${\bm{U}}^{r}$. Thus, we can mix the features between blocks without permuting the inner features. The formula is defined as ${\bm{W}}{\bm{x}}\approx{\bm{U}}^{r}({\bm{V}}^{b}{\bm{x}})$, where ${\bm{V}}^{b}$ is the block-diagonal matrix with $b$ blocks in shape $\frac{R}{B}\times\frac{N}{B}$, and ${\bm{U}}^{r}\in\mathbb{R}^{M\times R}$. Technically, the second projection does not need to be a low-rank approximation because $R$ can be larger than $M$. Nevertheless, in practice, we chose $R<M$ to limit the search space of this work, and thus use the superscript $r$ for the second matrix. The parameters of this method are determined by two variables $B$ and $R$, cutting the original burden from $M\cdot N$ to $R\cdot(M+\frac{N}{B})$. Note that BlockDense can recover the LowRank approximation if we set $B=1$ and $R<\min(M,N)$.
Remark
We limit our exploration of the efficient linear parametrizations within the FFN blocks.These typically have $8\cdot H^{2}$ parameters and MAC, where $H$ standards for hidden state dimension and $4\cdot H$ as the intermediate hidden size of FFN. In contrast, the proposed parametrizations have:
@@
Although BlockDense is introduced as a new parameterization, the aim of this paper is not to claim it as the best candidate, but rather to investigate some general properties of structured matrices from efficiency, optimization, and scaling perspectives. Given the favorable efficiency and loss performance of BlockDense, it is included alongside LowRank and BlockShuffle here to cover a broader range of potential parameterizations.
@@ 2.3 Addressing the optimization challenge
Using the efficient parametrization from initialization can suffer from optimization difficulty because the deep linear parametrization introduces additional symmetries [^2] Symmetries arise because such a factorization ${\bm{U}}{\bm{V}}$ is not unique; for any invertible matrix ${\bm{C}}$ of size $R\times R$ we have ${\bm{U}}{\bm{V}}={\bm{U}}{\bm{C}}{\bm{C}}^{-1}{\bm{V}}=({\bm{U}}{\bm{C}})({\bm{C}}^{-1}{\bm{V}})=\tilde{{\bm{U}}}\tilde{{\bm{V}}}$. BlockShuffle can also be included by fusing the shuffle operation into block-diagonal matrices. , which is a source of proliferation of saddle points and generally less smooth loss function as pointed out in [18, 19]. We hypothesize that this makes poorer learning dynamics of the structured parametrization. Empirically, we found that the deep linear form ${\bm{U}}({\bm{V}}{\bm{x}})$ is more difficult to train than the standard linear layer. For example, in Fig. 3, it can suffer from training instability and loss spikes with a large learning rate, while also converging much more slowly than the dense form with a small learning rate. We further elaborate on this, highlighting how the inconsistency of gradient updates between the structured parametrization and original linear projection affect learning dynamics in Appendix B.2.
@@
(b) Slow convergence
Figure 3: Poor training dynamics: Training dynamics of LowRank with rank of 128 under different training configurations. Curves correspond to a 4-layer Transformer with a model width of 768 on WikiText-103. We apply self-guided training in the first half of training. Refer to Sec. B.1 for more training dynamics visualizations of the other two structured parameterizations.
@@
To motivate our proposal, we start by finding that the updates on ${\bm{U}}{\bm{V}}$ scales the function of the backpropagated gradients ${\bm{g}}$ (see App. B.2), then turn into the typical training dynamics with gradients. An issue that needs to be addressed in the early stage of training is feature specialization when the learning process assigns semantics to the different hidden units of the model, sometimes also phrased as identifying the winning ticket [20]. In this process, certain weights will need to be suppressed, and symmetries in the parametrization must be resolved.
@@
Guided by the dense weights, which do not have the symmetry problem, it becomes much easier for the structured matrices to learn a good representation. From Fig. 3, it can be observed that the self-guided training prevents training spikes and fastens the convergence process. Notably, it benefits all three structured matrices with improved training dynamics illustrated in Sec. B.1 and better final performance shown in Sec. 4.4.
@@ 3 Related work
Efficient techniques for training LLMs Recent advancements in attention mechanisms have significant improvements in the efficiency of attention [21, 22, 23, 2, 24, 25, 26, 27, 28, 29], and the focus has shifted towards improving FFNs in LLMs, which contributes to at least half of the training time. Dynamic architectures such as mixtures of experts [30, 31, 32], or optimizers with faster convergence [33, 34] have been popular in improving training efficiency. Moreover, Xia et al. [35] employs structured pruning with learned sparsity masks and a dedicated data-loading policy to reduce the training budget.
@@ 4 Experiments
In our experiments, we empirically analyzed the performance of scaling, efficiency, and self-guided training for structured parameterization in LLMs.
4.1 Settings
Model
We perform the experiments on the standard Transformer architecture [45, 4] equipped with rotary positional embeddings [46] and the Llama Tokenizer [47]. Its FFN module is composed of two linear layers and a GeLU activation. Four sizes are considered, including Transformer-s (110M), Transformer-m (335M), Transformer-l (729M), and Transformer-xl (1.3B). For our efficient parameterizations, we only make the FFN module structured in most experiments to simplify our study, as the attention module has been well-studied [2, 23]. We explore two sizes that retain 63% or 32% of the dense FFN parameters by adjusting the rank and number of blocks (e.g., using a rank half the FFN width in LowRank reduces the parameters to 63%). In particular, to provide more comparative results with dense models in the scaling study Sec. 4.2, we further design the wide and structured networks, where both the attention and FFN modules are made structured using  [2] and the structured matrices. This is because allocating more parameters to the FFN compared to the attention module is more favorable, and making them both structured helps maintain the parameter ratio between them. Detailed configurations can be found in Table 10.
For implementation, we take Dao et al. [13]’s implementation for the BlockShuffle and carefully manage memory copies for BlockDense. In our experiments, we chose $B$ as a common divisor of $M$ and $N$ or $R$. Proper initialization is also investigated in Sec. A.1.
Training
@@ 4.1 Settings
We use the RefinedWeb dataset [48] and randomly select 0.5B tokens for validation, reserving the rest for training. All experiments, except for the overtraining experiments on 300B tokens in Fig. 5, are based on the Chinchilla scaling law [1], where tokens are allocated at 20 times the baseline model size. We set hyperparameters such as learning rates and global batch size according to the scaling law studies from recent papers [49, 50]. However, for the 300B token experiments, we found that more advanced hyperparameter settings are necessary. For example, we use betas of [0.9, 0.98]. Additionally, different works tend to use very different learning rates [51, 49, 52, 47] in the overtraining regime. Thus, we follow the scaling law of hyperparameters described in Bi et al. [53] to avoid extensive hyperparameter searches. Other implementations include using A100 80G GPUs, mixed precision (bfloat16 and float32), and adopting fast CUDA kernels like Flash Attention [25] for all experiments. We measure training FLOPs as in Megatron [54], including all matrix multiplications. Additional details are provided in Appendix C.
4.2 Scaling analysis
We evaluate the scaling performance of structured linear parameterizations from two perspectives. The first study investigates the scaling law of training-time compute. The second study trains these models with 300B tokens and evaluates their performance on downstream tasks. The results show that our structured matrices can serve as a strong alternative to the dense FFN, utilizing training FLOPs more efficiently (e.g., smaller model and lower loss) and performing better in the overtraining regime.
Scaling law study: better training FLOPs utilization